### v1

# Structure of the mean-field yrast spectrum of a two-component Bose gas in a ring: role of interaction asymmetry

Abstract
The mean-field yrast spectrum of an SU(2)-symmetric two-component Bose gas confined to a ring geometry is known to exhibit an intricate nonanalytic structure that is absent in single-component systems. In particular, due to the interplay between the species concentration and the atomic interactions, a sequence of plane-wave states can emerge as yrast states at fractional values of the angular momentum per particle. This behavior stands in sharp contrast to the single-component case, where plane-wave states occur only at integer angular momenta. In this paper, we investigate how the structure of the yrast spectrum in a two-component Bose gas is modified by interaction asymmetry. By numerically solving the coupled Gross–Pitaevskii equations for propagating soliton states, we compute the mean-field yrast spectrum and, in particular, determine the critical curves associated with the emergence of various plane-wave yrast states. We find that both the behavior of these critical curves and the mechanisms by which plane-wave yrast states arise depend sensitively on the relative strengths of the inter- and intra-component interactions. When the inter-component interaction is weaker, the plane-wave yrast states replace soliton states through a continuous evolution, as in the SU(2)-symmetric case, although the conditions for their existence become more restrictive. In contrast, when the inter-component interaction is stronger, plane-wave yrast states emerge by overtaking soliton states via branch crossings, and their stability is significantly enhanced. Our results have important implications for the existence and stability of persistent currents in asymmetric, two-component Bose gases.
I Introduction
The study of persistent currents in neutral superfluids confined to a multiply connected region originated from efforts to understand analogous phenomena in superconducting rings. Using arguments similar to those of Byers and Yang in their explanation of magnetic flux quantization in superconducting rings [Yang61], Bloch showed that the occurrence of persistent currents in neutral superfluids in a ring arises from the characteristic dependence of the superfluid energy on its angular momentum per particle [Bloch73]. This energy spectrum, sometimes referred to as the yrast spectrum [Mottelson99], takes the general form
$E_{0}(l)=\frac{l^{2}\hbar^{2}}{2MR^{2}}+e_{0}(l),$ (1)
where $E_{0}(l)$ is the superfluid energy per particle, $M$ is the mass of the atom, $R$ is the radius of the ring, and $l\hbar$ is the angular momentum per particle. Here $e_{0}(l)$ represents the internal energy which has inversion symmetry $e_{0}(l)=e_{0}(-l)$ and possesses the periodicity property $e_{0}(l+n)=e_{0}(l)$ where $n$ is an integer (see Fig. 1). These properties, along with the general assumption that $e_{0}(l)$ has a finite slope as $l$ approaches zero, imply that the yrast spectrum is not analytic at integer values of angular momentum per particle. Under suitable conditions these non-analytic points can emerge as local minima of the yrast spectrum which support persistent currents. Characteristically, the circulation of the fluid is quantized at these integer values of angular momentum per particle. This is analogous to the situation in a superconducting ring, where persistent currents are tied to the quantization of magnetic flux [Deaver61, Doll61].
Figure 1: Illustration of the yrast spectrum for a single-component superfluid in a ring of radius $R$.
In recent decades, quantum gases have become as an important platform for investigating persistent currents for two principal reasons [Ryu07, Ramanathan11, Moulder12, Beattie13, Wright13, Cai22, Del22, Polo25]. First, experimental advances in realizing multicomponent superfluids and synthetic gauge fields have substantially broadened the class of systems that may support persistent currents [KAWAGUCHI12, Lin_2016]. Second, the yrast spectrum of such systems can be determined quantitatively by mean-field theory, revealing a rich structure and enabling persistent currents beyond the conventional single-component scenario [Smyrnakis09, Anoshkin2013, Wu2013, Abad14, Wu2015, Pecci23]. Indeed, the mean-field yrast spectrum of a SU(2)-symmetric two-component Bose gas in a ring geometry is known to exhibit an intricate structure not present in single-component systems [Smyrnakis09, Anoshkin2013, Wu2013]. Although it still obeys the general form given in Eq. (1), the yrast spectrum of this two-component system can develop a series of non-analytic points at certain fractional values of $l$, arising from the interplay between component concentration and atomic interactions. The states at these additional points, much like those at integer $l$, may also support persistent currents. This feature of the yrast spectrum was first uncovered using analytic soliton solutions of the coupled Gross–Pitaevskii equations for the SU(2)-symmetric system, and its implications for persistent currents were later confirmed experimentally [Beattie13]. Notably, the appearance of each additional non-analytic point at the fractional $l$ is accompanied by a transition in the condensate wave functions: a pair of soliton solutions is replaced by a pair of plane-wave states. Moreover, these plane-wave states carry different winding numbers, reflecting distinct quantizations of circulation in the two components.
Most experimentally accessible two-component systems are not SU(2) symmetric, since the intra- and inter-component interaction strengths generally differ. In this case, the coupled Gross–Pitaevskii (GP) equations no longer admit analytic solutions, and it is therefore natural to ask whether the recently discovered features in the yrast spectrum persist under these more realistic conditions. Reference [Wu2015] attempted to address this question using a perturbative approach based on an assumption motivated by the SU(2)-symmetric case—namely, that the appearance of a nonanalytic point at a fractional value of the angular momentum $l$ is always accompanied by a continuous transformation of the corresponding soliton state into a plane-wave state at that angular momentum. However, this assumption has never been verified for asymmetric systems. Several other works [Smyrnakis2014, Sandin16, Matsushita2018, Roussou2018] have also numerically investigated the yrast spectrum of asymmetric two-component systems, but an overarching picture of the structure of the yrast spectrum has yet to emerge. In this work, we present a comprehensive study of the mean-field yrast spectrum of asymmetric two-component Bose gases confined to a ring geometry by numerically solving the coupled GP equations. We find that the assumption adopted in Ref. [Wu2015] holds only for asymmetric two-component gases in which the inter-component interaction is weaker than the intra-component interaction. In the opposite regime, this assumption breaks down: the plane-wave yrast states instead emerge from crossings between distinct branches of solutions to the GP equations, and the resulting yrast spectrum exhibits an even richer nonanalytic structure than in the SU(2)-symmetric case. In both regimes, we establish critical conditions for the emergence of plane-wave yrast states, which to a large extent determine the analytic structure of the yrast spectrum.
The rest of the paper is organized as follows. In Sec. II, we review the known analytic properties of the yrast spectrum of an SU(2)-symmetric system, with particular emphasis on the critical conditions for the emergence of plane-wave yrast states. The main objective of this paper is then explicitly stated in this section: to determine the corresponding critical conditions for the asymmetric system through numerical solutions of the coupled Gross–Pitaevskii equations. In Sec. III, we develop an efficient numerical method used in solving these equations and present examples of the yrast spectrum calculated using the solutions. The structure of the yrast spectrum can be understood by the critical conditions and phase diagrams of the plane-wave yrast states which are presented in Sec. IV. The main results are summarized in Sec. V.
II Statement of the problem
We consider a two-component Bose gas of $N$ atoms confined to a ring of radius $R$, where $N_{A}$ atoms are in hyperfine state $A$ and $N_{B}$ atoms in hyperfine state $B$. The atoms interact via contact interactions of strengths $U_{ss^{\prime}}$ where $s$ and $s^{\prime}$ label the components $A$ and $B$. We assume that $N_{B}<N_{A}$, so that $B$ is by definition the minority component, and that the two components have the same mass $M$. Within a mean-field Gross-Pitaevskii (GP) description, the yrast spectrum (in units of $N\hbar^{2}/2MR^{2}$) can be calculated by
$\bar{E}_{0}(l)=$ $\sum_{s}x_{s}\int_{0}^{2\pi}d\theta\left|\psi^{\prime}_{s}(\theta)\right|^{2}$
$+$ $\pi\sum_{ss^{\prime}}x_{s}x_{s^{\prime}}\gamma_{ss^{\prime}}\int_{0}^{2\pi}d\theta|\psi_{s}(\theta)|^{2}|\psi_{s^{\prime}}(\theta)|^{2},$ (2)
where $x_{s}=N_{s}/N$ is the component concentration and $\gamma_{ss^{\prime}}=U_{ss^{\prime}}NMR^{2}/\pi\hbar^{2}$ are three dimensionless interaction parameters. The condensate wave functions $\psi_{A}(\theta)$ and $\psi_{B}(\theta)$ are the lowest energy solutions to the coupled time-independent Gross-Pitaevskii equations
$-\psi^{\prime\prime}_{s}+i\Omega\psi^{\prime}_{s}+2\pi\sum_{s^{\prime}}\gamma_{ss^{\prime}}x_{s^{\prime}}|\psi_{s^{\prime}}|^{2}\psi_{s}=\mu_{s}\psi_{s}$ (3)
at a fixed angular momentum per particle $l\hbar$. Here $\mu_{s}$ and $\Omega$ are (dimensionless) Lagrange multipliers associated with the normalization constraints
$\int_{0}^{2\pi}d\theta|\psi_{s}(\theta)|^{2}=1,$ (4)
and the angular momentum constraint
$\frac{1}{i}\sum_{s}x_{s}\int_{0}^{2\pi}d\theta\psi_{s}^{*}(\theta)\psi^{\prime}_{s}(\theta)=l.$ (5)
In general, the equations in (3) in fact describe solitons that travel at the angular velocity $\Omega$ in the ring. Although the three interaction parameters may, in principle, all differ, we focus on the experimentally relevant case in which the intra-component interactions are equal but differ from the inter-component interaction, i.e., $\gamma_{AA}=\gamma_{BB}\neq\gamma_{AB}$. We therefore set
$\gamma_{AB}$ $=\gamma;$
$\gamma_{AA}$ $=\gamma_{BB}=(1+\kappa)\gamma,$ (6)
where the dimensionless parameter $\kappa$ characterizes the interaction asymmetry. In the special case of $\kappa=0$, the system possesses the SU(2) symmetry and Eq. (3) can be solved analytically in terms of the Jacobi elliptic functions.
Figure 2: The internal part of the mean-field yrast spectrum of an SU(2)-symmetric system, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=100$. The circles indicate the locations at which the derivative of the spectrum is discontinuous.
The purpose of our paper is to determine the analytic structure of the mean-field yrast spectrum $\bar{E}_{0}(l)$ as a function of $l$ for the asymmetric system. To phrase our objective more precisely, we first review the known properties of the yrast spectrum and the associated condensate wave functions in the SU(2)-symmetric system. Because of the general form in Eq. (1), we need only to consider the internal energy (in units of $N\hbar^{2}/2MR^{2}$)
$\bar{e}_{0}(l)\equiv\bar{E}_{0}(l)-l^{2}$ (7)
in the fundamental range $0<l\leq 1/2$. Unlike the single-component gas, for which the yrast spectrum in this range is a smooth curve, the derivative of the spectrum in the SU(2)-symmetric two-component system exhibits discontinuities at $l=kx_{B}$, where k=1,2,,Kk=1,2,...,K (see Fig. 2). A derivative discontinuity at a given value $l=kx_{B}$ appears when the parameters $(\gamma,x_{B})$ lie within the region of the $\gamma$-$x_{B}$ plane bounded by the corresponding critical curve $x_{B}(\gamma,q)$ and the $x_{B}=0$ axis. The critical curve for $k=1$ is simply the line $x_{B}=1/2$, implying that the discontinuity at $l=x_{B}$ is always present. The critical curves for $k=2,3,4$ are illustrated by the solid lines in Fig. 3; each approaches the asymptotic value $\big(1-\sqrt{1-1/k^{2}}\big)/2$ as $\gamma\rightarrow\infty$. Thus, the spectrum contains a total $K$ of derivative discontinuities when the coordinate $(\gamma,x_{B})$ lies within a region bounded by the two critical curves $x_{B}(\gamma,K)$ and $x_{B}(\gamma,K+1)$ [Wu2015], as shown in Fig. 3. In light of this diagram, the variation of the yrast-spectrum structure with $x_{B}$, shown in Fig. 2, is readily understood.
Figure 3: Critical curves for plane-wave yrast states $(\phi_{0},\phi_{k})$ with $k=2,3,4$ and the corresponding phase diagram for the SU(2)-symmetric system. Note that the $\gamma$ axis is plotted on a logarithmic scale.
The non-analyticity at these points indicates that the corresponding condensate wave functions are special compared with those at generic values of $l$. Indeed, at the non-analytic $l=kx_{B}$ the condensate wave function takes a plane-wave form $(\psi_{A},\psi_{B})=(\phi_{0},\phi_{k})$ where
$\phi_{k}=e^{ik\theta}/\sqrt{2\pi},$ (8)
whereas the yrast state is generally a soliton state at other fractional values of $l$. Consequently, the critical curves $x_{B}(\gamma,k)$ in Fig. 3 can also be interpreted as delineating the phases in which different numbers of plane-wave yrast states appear. It is interesting to see how these plane-wave yrast states emerge from soliton states as the interaction strength changes. For example, consider a system with $x_{B}=0.02$ and increase $\gamma$ from a small value (see the dashed line with an arrow in Fig. ). Since this value of $x_{B}$ lies within the range
$\frac{1}{2}\left[1-\sqrt{1-1/(k+1)^{2}}\right]<x_{B}<\frac{1}{2}\left[1-\sqrt{1-1/k^{2}}\right]$
with $k=3$, the yrast spectrum has at most three derivative discontinuities in the fundamental range. As noted earlier, the yrast state at $l=x_{B}$ is always $(\phi_{0},\phi_{1})$, independent of $\gamma$. The yrast spectrum develops a second derivative discontinuity at $l=2x_{B}$ once $\gamma>9$, a value determined by the intersection of the line $x_{B}=0.02$ with the critical curve $x_{B}(\gamma,2)$. As $\gamma$ approaches this critical value from below, the soliton state at $l=2x_{B}$ continuously evolves into the plane-wave state $(\phi_{0},\phi_{2})$. When $\gamma$ reaches the critical value, this plane-wave state emerges and becomes pinned at $l=2x_{B}$ as $\gamma$ increases further. Finally, a third and final plane-wave yrast state $(\phi_{0},\phi_{3})$ emerges at $l=3x_{B}$ when $\gamma>60$, as determined by the intersection of the same line with the critical curve $x_{B}(\gamma,3)$.
The above summary shows that, for the SU(2)-symmetric system, the yrast spectrum in the fundamental range consists of smooth soliton branches joined by plane-wave states $(\phi_{0},\phi_{k})$ at fractional values $l=kx_{B}$, where each junction appears as a non-analytic point. Owning to the periodicity and the inversion symmetry of the internal energy, the appearance of the plane-wave yrast state $(\phi_{0},\phi_{k})$ necessarily implies the simultaneous appearance of the states $(\phi_{\mu},\phi_{\mu\pm k})$ where $\mu=1,2,\cdots$. In other words, the analytic structure of the yrast spectrum is fundamentally determined by the emergence of these plane-wave yrast states. We can therefore state our problem more concisely: for the asymmetric system, which plane-wave states can become the yrast states, and under what conditions? Or, equivalently, what are the critical curves and phase diagram of the plane-wave yrast states for the asymmetric system? In addressing this question using numerical solutions of the GP equations, we will also determine whether the emergence of plane-wave yrast states from soliton states always occurs in a continuous manner.
III Numerical method
We employ the method of imaginary time propagation to obtain solutions to the coupled GP equations in Eq. (3) for $0\leq l\leq 1/2$, which are then substituted in Eq. (2) to calculate the yrast spectrum. Our method differs from that employed in Ref. [Sandin16]. The key challenge in solving Eq. (3) is to implement the angular momentum constraint in Eq. (5), or equivalently, to determine the Lagrange multiplier $\Omega$ in terms of $l$ and the condensate wave functions. For this purpose, we make use of the modulus-phase representation
Figure 4: The internal part of the mean-field yrast spectrum of an asymmetric system with $\kappa=-0.1$, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=20$. The circles indicate the locations at which the derivative of the spectrum is discontinuous.
$\psi_{s}(\theta)=\sqrt{\rho_{s}(\theta)}e^{i\varphi_{s}(\theta)}.$ (9)
Figure 5: The internal part of the mean-field yrast spectrum of an asymmetric system with $\kappa=1$, plotted in the fundamental range $0<l\leq 1/2$ for $\gamma=20$. The circles indicate the locations at which the derivative of the spectrum is discontinuous.
The single-valuedness of the wave function $\psi_{s}(\theta)$ implies the following boundary conditions
$\rho_{s}(\theta+2\pi)-\rho_{s}(\theta)$ $=0$ (10)
$\varphi_{s}(\theta+2\pi)-\varphi_{s}(\theta)$ $=2\pi J_{s},\quad J_{s}=0,\pm 1,\pm 2,\cdots$ (11)
where the integers $J_{s}$ are referred to as phase winding numbers. Substituting Eq. (9) into Eq. (3), one obtains
$\frac{\rho^{\prime\prime}_{s}}{2\rho_{s}}+\frac{\rho^{\prime}_{s}}{4\rho^{2}_{s}}+\left(\varphi^{\prime}_{s}-\Omega\right)\varphi^{\prime}_{s}+2\pi\sum_{s^{\prime}}\gamma_{ss^{\prime}}x_{s^{\prime}}\rho_{s^{\prime}}=\mu_{s}$ (12)
and
$\sqrt{\rho_{s}}\varphi^{\prime\prime}_{s}+\left(2\varphi^{\prime}_{s}-\Omega\right)\left(\sqrt{\rho_{s}}\right)^{\prime}=0.$ (13)
Equation. (13) can be solved as
$\varphi^{\prime}_{s}(\theta)=\frac{W_{s}}{2\rho_{s}}+\frac{\Omega}{2},$ (14)
where $W_{s}$ is an integration constant. Integrating the above equation and using Eq. (11), we find
$W_{s}=\frac{4\pi J_{s}-2\pi\Omega}{I_{s}},$ (15)
where
$I_{s}\equiv\int_{0}^{2\pi}d\theta\frac{1}{\rho_{s}(\theta)}.$ (16)
Using Eq. (5) and Eq. (14), we find that the angular momentum per particle can be written as
$l=x_{A}l_{A}+x_{B}l_{B},$ (17)
where
$l_{s}={\pi W_{s}}+\frac{\Omega}{2}$ (18)
is the angular momentum per particle of each species. Substituting Eq. (15) in Eq. (17) we finally arrive at
$\Omega=\frac{l-4\pi^{2}(x_{A}J_{A}/I_{A}+x_{B}J_{B}/I_{B})}{1/2-2\pi^{2}(x_{A}/I_{A}+x_{B}/I_{B})}.$ (19)
We have thus expressed $\Omega$ in terms of $l$, as well as the modulus and the phase winding numbers of the condensate wave functions. In the imaginary time propagation of Eq. (3) at a specific $l$, the wave functions obtained after each iteration are used to evaluate $\Omega$ through the above expression, and the resulting $\Omega$ is then fed back into the next iteration. This leads to a very efficient algorithm for solving the coupled GP equations. As a benchmark, our numerical solutions for the SU(2)-symmetric case accurately reproduce the yrast spectrum obtained from the analytic soliton solutions, as well as the critical curves and phase diagram shown in Fig. 3.
Figure 6: The critical curves for the plane-wave yrast state $(\phi_{0},\phi_{2})$ at $\kappa=-0.1$ (solid line) and $\kappa=0$ (dashed line).
Figure 7: Critical curves for the emergence of plane-wave yrast states and the corresponding phase diagram for $\kappa=-0.1$. Here value of $K$ indicates the total number of plane-wave yrast states in that phase.
In Fig. 4, we present examples of the internal yrast spectrum calculated for several values of $x_{B}$ at $\kappa=-0.1$ and $\gamma=20$. We observe that the spectrum exhibits no derivative discontinuity for $x_{B}=0.1$, a single discontinuity at $l=0.05$ for $x_{B}=0.05$, and two discontinuities at $l=0.02$ and $l=0.04$ for $x_{B}=0.02$. Similar results for $\kappa=1$ and $\gamma=20$ are shown in Fig. 5, where the spectrum exhibits one derivative discontinuity at $l=0.4$ for $x_{B}=0.4$, two discontinuities at $l=0.2$ and $l=0.4$ for $x_{B}=0.2$, and three discontinuities at $l=0.1$, $0.2$, and $0.3$ for $x_{B}=0.1$. At all of these discontinuities, located at $l=kx_{B}$, the condensate wave functions correspond to the plane-wave states $(\phi_{0},\phi_{k})$. All of these features of the yrast spectrum will become transparent once the critical conditions for the emergence of plane-wave yrast states, discussed below, are established.
IV Critical conditions and phase diagrams
Figure 8: The upper and middle panels, respectively, show the amplitude and phase of the condensate wave function at $l=2x_{B}$ with $x_{B}=0.02$ for several values of $\gamma$. The continuous transition from the soliton state to the plane-wave state $(\phi_{0},\phi_{2})$ is most clearly illustrated by a plot of the amplitude as a function of the phase in the polar coordinate, displayed in the bottom panel.
It is clear that the critical curve $x_{B}(\gamma,k)$ for the emergence of the $(\phi_{0},\phi_{k})$ yrast state depends on the parameter $\kappa$ characterizing the interaction asymmetry. Because of the inversion symmetry we can restrict the possible $k$ to positive integers. We now present a simple argument showing that the behavior of these critical curves differ qualitatively for $\kappa<0$ and $\kappa>0$. From Eq. (2) we find that the energy of the plane-wave state $(\phi_{0},\phi_{k})$ is given by
$\bar{E}_{0}=x_{B}k^{2}+\frac{1}{2}\left[\gamma+\kappa\gamma(x_{A}^{2}+x_{B}^{2})\right].$ (20)
The energy difference $\delta\bar{E}_{0}$ between an arbitrary state $(\psi_{A},\psi_{B})$ in the same $l=kx_{B}$ manifold and this plane-wave state naturally separates into kinetic and interaction contributions, i.e.,
$\delta\bar{E}_{0}(\kappa)=\delta\bar{E}_{\rm kin}+\delta\bar{E}_{\rm int}(\kappa),$ (21)
where the kinetic energy difference does not depend on $\kappa$ and the interaction energy difference is given explicitly by
$\delta\bar{E}_{\rm int}(\kappa)$ $=\pi\kappa\gamma\left[x_{A}^{2}\int_{0}^{2\pi}d\theta|\delta\rho_{A}|^{2}+x_{B}^{2}\int_{0}^{2\pi}d\theta|\delta\rho_{B}|^{2}\right]$
$+\pi\gamma\int_{0}^{2\pi}d\theta|\delta\rho|^{2}$ (22)
with $\delta\rho_{s}(\theta)=\rho_{s}(\theta)-1/2\pi$ and $\delta\rho(\theta)=x_{A}\delta\rho_{A}(\theta)+x_{B}\delta\rho_{B}(\theta)$. Now suppose that the plane-wave state $(\phi_{0},\phi_{k})$ is an yrast state at $\kappa=0$ for some parameter set $(\gamma,x_{B})$, implying that it has the lowest energy in the $l=kx_{B}$ manifold, i.e.,
$\delta\bar{E}_{0}(\kappa=0)=\delta\bar{E}_{\rm kin}+\delta\bar{E}_{\rm int}(\kappa=0)>0.$ (23)
Under the same parameters, this plane-wave state is then guaranteed to remain an yrast state for $\kappa>0$, since
$\delta\bar{E}_{0}(\kappa>0)\geq\delta\bar{E}_{0}(\kappa=0).$ (24)
Conversely, this state may cease to be an yrast state when $\kappa<0$, since
$\delta\bar{E}_{0}(\kappa<0)\leq\delta\bar{E}_{0}(\kappa=0).$ (25)
The above analysis shows that an asymmetric interaction with $\kappa>0$ stabilizes the plane-wave state as an yrast state, whereas $\kappa<0$ tends to destabilize it. Given the saturating behavior of the critical curves at $\kappa=0$ shown in Fig. , we expect the curves for $\kappa<0$ to tilt downward and move away from the asymptotic lines, while those for $\kappa>0$ tilt upward and cross these lines.
IV.1 $\kappa<0$
As an example, we show in Fig. 6 the critical curve for the plane-wave yrast state $(\phi_{0},\phi_{2})$ at $\kappa=-0.1$, alongside the corresponding curves for $\kappa=0$. The curve at $\kappa=-0.1$ behaves exactly as anticipated from the preceding argument. In particular, the critical values of $x_{B}$ all approach zero at large $\gamma$. Since the plane-wave state $(\phi_{0},\phi_{k})$ becomes an yrast state in the region bounded by its critical curve and the $x_{B}=0$ axis, this implies that the state will eventually cease to be an yrast state at sufficiently large $\gamma$, regardless of the value of $x_{B}$. Such a behavior can also be understood from Eqs. (21) and (22), which show that $\delta\bar{E}_{0}(\kappa<0)$ decreases without bound as $\gamma$ increases and will ultimately become negative. In other words, for a fixed negative value of $\kappa$, increasing $\gamma$ also tends to destabilize the plane-wave state as an yrast state. This is most clearly illustrated by the $(\phi_{0},\phi_{1})$ state: it is always an yrast state at $\kappa=0$, yet for $\kappa<0$ it ceases to be one once $\gamma$ exceeds a certain threshold (see Fig. 7). In fact, the entire system becomes unstable once $\gamma$ exceeds values set by the dynamical stability condition
$\left[(1+\kappa)\gamma+{1}/{2}\right]^{2}>\gamma^{2}.$ (26)
The stability curve together with a partial phase diagram of the plane-wave yrast states is shown in Fig. 7 for $\kappa=-0.1$. A complete phase diagram is not feasible to present, as phases hosting an increasing number of plane-wave yrast states emerge at progressively smaller values of $x_{B}$ and increasingly larger values of $\gamma$.
Interestingly, we find that the critical curves for $\kappa<0$ obtained from numerical solutions of the coupled GP equations agree with those determined using a perturbative approach. As discussed in Sec. I, this implies that the assumption of a continuous transition between the soliton states and the plane-wave yrast states at $l=kx_{B}$ is correct in this case. This is verified by our numerical calculations. As a representative example, we fix $x_{B}=0.02$ and increase $\gamma$. As shown in the phase diagram in Fig. 6, the plane-wave state $(\phi_{0},\phi_{2})$ becomes an yrast state once $\gamma$ exceeds the critical value $\gamma_{c}\simeq 11.2$. In Fig. 8, we plot the condensate wave functions at $l=2x_{B}$ for $\gamma=8$, $10$, $11$, and $11.5$. These results clearly demonstrate a continuous evolution from soliton states to the plane-wave state $(\phi_{0},\phi_{2})$ as $\gamma$ approaches and crosses the critical value.
IV.2 $\kappa>0$
Since $(\phi_{0},\phi_{1})$ is always an yrast state for $\kappa=0$, the preceding analysis implies that it remains an yrast state for $\kappa>0$ as well. Turning to the critical curve of the plane-wave state $(\phi_{0},\phi_{2})$ for positive $\kappa$, we find that it indeed crosses the asymptotic line at $\kappa=0$ as $\gamma$ increases, as illustrated in Fig. 9. Interestingly, additional structure emerges beyond this crossing. Once the asymptotic line at $\kappa=0$ is crossed, the critical value of $x_{B}$ continues to increase with $\gamma$ and eventually saturates at $x_{B}=1/3$. However, between this curve and the line $x_{B}=1/2$, a lobe-shaped region develops within which $(\phi_{0},\phi_{2})$ is also an yrast state. Moreover, the upper and lower boundaries of this lobe approach $x_{B}=1/2$ and $x_{B}=1/3$, respectively, in the asymptotic limit. In other words, the horizontal lines $x_{B}=1/3$ and $x_{B}=1/2$ act as asymptotes for the $k=2$ critical curve. We therefore conclude that for $\kappa>0$, the plane-wave state $(\phi_{0},\phi_{2})$ eventually becomes an yrast state at sufficiently large $\gamma$ for all values of $x_{B}$, except precisely at $x_{B}=1/3$ and $x_{B}=1/2$.
Figure 9: The critical curves for the plane-wave yrast state $(\phi_{0},\phi_{2})$ at $\kappa=0$ (solid line) and $\kappa=1$ (dashed line).
Figure 10: Critical curves for plane-wave yrast states $(\phi_{0},\phi_{k})$ with $k=2,3$ and the corresponding phase diagram for $\kappa=1$.
To understand this phenomenon, we revisit the analysis at the beginning of this section, where the energies of the plane-wave state $(\phi_{0},\phi_{k})$ and another state $(\psi_{A},\psi_{B})$ within the same $l=kx_{B}$ manifold are compared. If $(\psi_{A},\psi_{B})$ is a soliton state for which $|\delta\rho_{A}|\neq 0$ and $|\delta\rho_{B}|\neq 0$, the first term of $\delta\bar{E}_{\rm int}(\kappa>0)$ in Eq. (22) can be made arbitrarily large by increasing $\gamma$. Consequently, even if $(\phi_{0},\phi_{k})$ is not an yrast state for $\kappa=0$, its energy at $\kappa>0$ can become lower than that of any soliton state in the $l=kx_{B}$ manifold for sufficiently large $\gamma$, irrespective of the value of $x_{B}$. However, this observation alone does not guarantee that $(\phi_{0},\phi_{k})$ will become an yrast state, since there may exist other plane-wave states within the same $l=kx_{B}$ manifold with even lower energies. To assess this possibility, we note that the minority concentration can in general be expressed as
$x_{B}=\frac{p}{q},$ (27)
where $p$ and $q$ are coprime integers and $p\leq 2q$. In the $l=kx_{B}$ manifold, there exist infinitely many plane-wave states of the form
$(\psi_{A},\psi_{B})=(\phi_{mp},\phi_{k-mq}),$ (28)
with $m=0,\pm 1,\pm 2,\ldots$, all of which share the same interaction energy. These states are distinguished by their kinetic energies,
$\bar{E}_{\rm kin}=l^{2}+x_{A}x_{B}(k-mq)^{2},$ (29)
and only the plane-wave state with the lowest kinetic energy can potentially serve as the yrast state. Since $k$ is restricted to positive integers, this condition implies that for $(\phi_{0},\phi_{k})$ to be a viable candidate for the yrast state, $k$ must lie within the range
$k\leq\left\lfloor\frac{q}{2}\right\rfloor,$ (30)
where $\lfloor q/2\rfloor$ denotes the integer part (floor) of $q/2$. Once $k$ lies within this range, then the previous argument implies that $(\phi_{0},\phi_{k})$ will eventually become an yrast state for sufficiently large $\gamma$. Consequently, for a fixed $x_{B}=p/q$, the plane-wave states $(\phi_{0},\phi_{k})$ emerge as yrast states one by one with increasing $\gamma$, in order of increasing $k$, up to $k=\lfloor q/2\rfloor$. Equation  30 also means that the plane-wave state $(\phi_{0},\phi_{k})$ can never become an yrast state for those concentrations $x_{B}=p/q$ satisfying $\lfloor q/2\rfloor<k$, regardless of how large $\gamma$ is. This implies, for example, that $(\phi_{0},\phi_{2})$ can never become an yrast state at $x_{B}=1/2$ or $x_{B}=1/3$, which is precisely why these values appear as asymptotes of the critical curve $x_{B}(\gamma,k=2)$. Likewise, $(\phi_{0},\phi_{3})$ can never become an yrast state for $x_{B}=1/2,1/3,1/4,1/5,$ and $2/5$. This conclusion is confirmed by our numerical determination of the critical curve $x_{B}(\gamma,k=3)$, shown along with a partial phase diagram in Fig. 10, where these values indeed serve as asymptotes. The asymptotes of the critical curves for higher $k$ values can be deduced similarly.
Importantly, the critical curves determined here deviate substantially from those obtained perturbatively in Ref. [Wu2015]. This indicates that, for $\kappa>0$, plane-wave yrast states do not emerge through a continuous transformation from soliton states, as is the case for $\kappa\leq 0$. To substantiate this observation, we examine the evolution of the condensate wave functions at $l=2x_{B}$ for $x_{B}=0.2$ and $\kappa=1$ across the critical value $\gamma_{c}\simeq 5.2$, as indicated in Fig. 10. In Fig. 11, we plot the condensate wave functions at $\gamma=5$ and $\gamma=5.5$, corresponding to values just below and above the critical point. We observe that the condensate wave function at $\gamma=5$ is a soliton state that does not continuously evolve into the plane-wave state observed at $\gamma=5.5$. This clearly demonstrates that the emergence of the plane-wave yrast state occurs via a branch crossing rather than through a continuous transformation.
Figure 11: Plots of the amplitude as a function of the phase in the polar coordinate, displayed in the bottom panel.
V Conclusions
In this paper, we have systematically determined the critical curves governing the existence of plane-wave yrast states in an asymmetric two-component Bose gas confined to a ring, using numerical solutions of the coupled GP equations supplemented by analytic insights. Because the emergence of these plane-wave yrast states gives rise to derivative discontinuities in the yrast spectrum, the resulting critical curves define a phase diagram that provides a clear picture of the analytic structure of the yrast spectrum.
[…]
Acknowledgements
This work is supported by National Key R&\&D Program of China (Grant No. 2022YFA1404103), Natural Science Foundation of China (Grant No. 12474264), Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2404007). S.Z. acknowledges support from HK GRF (Grant No. 17306024), CRF (Grants No. C6009-20G, No. C7012-21G, No. C4050-23GF), and a RGC Fellowship Award No. HKU RFS2223-7S03.